# Imprints of the Local Bubble and Dust Complexity on Polarized Dust Emission

### v1

@@ 3.2 Testing the Dependence of the Dust Polarization Fraction on the Magnetic Inclination Angle
Using 3D dust extinction maps and a model of the Local Bubble surface geometry, we can measure the angle between the line of sight and the Bubble’s surface in each direction. If the magnetic field lines were tangential to the Local Bubble surface as discussed in Section 3.1, the angle between the line of sight and the Bubble’s surface should on average be correlated with the measured dust polarization fraction. Therefore, by quantifying the correlation between this angle and the dust polarization fraction for different sightlines, we can test whether there is measurable evidence that the magnetic field traced by the dust polarization is preferentially tangential to the Local Bubble surface.
@@
$Q~{}=\int\alpha\epsilon\rho\cos{2\psi}\cos^{2}{\gamma}\,ds,$ (3)
$U~{}=~{}\int\alpha\epsilon\rho\sin{2\psi}\cos^{2}{\gamma}\,ds,$ (4)
where $\rho$ is the volume density, $ds$ is a distance segment along the line of sight, $\epsilon$ is the dust emissivity, $\alpha$ is a coefficient defined in Equation 15 of Padoan et al. (2001) that is a product of polarization efficiency factors, such as the degree of dust alignment and the dust grain polarization cross-section, and $\psi$ is the angle between the projection of the magnetic field on the plane of the sky and North. These equations show the dependence of the polarization fraction on $\gamma$$\psi$$\alpha$, and $\epsilon$ and their variations along the line of sight. At large angular scales, the variation in the observed dust polarization is dominated by $\gamma$ and $\psi$ (Hensley et al., 2019; Planck Collaboration et al., 2020a).
@@ 3.3 Magnetic Fields in the Local Bubble Wall
    For the Local Bubble geometry, we use the 3D model developed by Pelgrims et al. (2020). Since this model is constructed based on the 3D dust map of Lallement et al. (2019), we use this dust map in this subsection. Example slices of the map with the model overlaid are shown at the top of Figure 2. The resolution of this dust map is 25 pc. We use a 3D Gaussian filter to smooth the data to that resolution. This smooths out spurious small-scale fluctuations in the data product that may affect the results.
@@
    $\alpha_{i}~{}=~{}\pi-\arccos{\left(\frac{\nabla A_{{V},i}^{\prime}\cdot\vec{r_{i}}}{|\nabla A_{{V},i}^{\prime}||\vec{r_{i}}|}\right)},$ (5)
@@
3.4 No Detected Imprint of the Local Bubble Wall on the Polarization Fraction
@@
We do not find any correlation between $p_{353}$ and $\theta$ from Equation 8 over the mask defined in Figure 3. The Spearman rank coefficient, which is agnostic to the functional dependence between $p_{353}$ and $\theta$, is $3~{}\times~{}10^{-3}$. This indicates that at least one of the assumptions described in Section 3.1 is not valid. In other words, either the magnetic field is not generally tangent to the Local Bubble surface, the dust polarization is not dominated by the dust within 50 pc of the surface as defined by the Pelgrims et al. (2020), or a combination of these possibilities.
@@
We also test whether we find an anti-correlation between $\theta$ and the absolute value of the rotation measure but do not find any evidence for it. Therefore, we do not find an imprint of the detailed 3D geometry of the Local Bubble wall on the polarization statistics.
@@
O’Neill et al. (2024) provide a map of the inclination angle between the Local Bubble wall as defined by their model and the plane of the sky. We also test for a correlation between this angle and the polarization fraction and do not find any.
4 Imprint of Dust Complexity on Polarization Fraction
@@
We continue the investigation of how the 3D distribution of dust impacts the polarization fraction beyond the Local Bubble, taking into account the distribution of the dust in an extended volume around the Sun. We use the 3D dust maps constructed by Edenhofer et al. (2023) in this section. The benefits of these maps are that they have high resolution and extend radially up to 1.25 kpc away from the Sun.
@@
If contributions to the polarized dust emission originate from regions along the line of sight with differently oriented magnetic fields, the integrated signal will be depolarized relative to emission from a region with uniform magnetic fields. We postulate that sightlines with multiple dust components that are separated in distance and contribute similarly to the total column density are more likely to have substantial dust emission originating from regions with differently oriented magnetic fields. We test the hypothesis that on average, for sightlines with the same column density, the ones for which the dust is distributed into multiple components at different distances with similar contributions to the total column density are associated with higher levels of depolarization than those for which the dust contribution to the total column density is concentrated.
@@ 4.1 Sightline Selection
Edenhofer et al. (2023) use estimates of stellar distances and extinctions from Zhang et al. (2023) to construct their 3D dust maps. However, Edenhofer et al. (2023) found that their predictions disagree with Zhang et al. (2023) where the radially integrated differential extinction is below 50 mmag or above 4 mag. We therefore use those thresholds to mask sightlines where the dust differential extinction is likely to be significantly over- or under-estimated. The distance to those thresholds for a given sightline varies slightly for the 12 different posterior samples, so we apply a slightly different mask to each posterior sample. We only consider sightlines where the differential extinction integrated radially outwards reaches 50 mmag within 200 pc of the Sun. A threshold higher than 200 pc would include more sightlines in the selection, but it would shorten the minimum path length considered. Because in Section 4.5, we compare our results with the Leike et al. (2020) maps, which only extend out to 270 pc, we find 200 pc to be a good balance between having a large enough sample size and minimum path length. The minimum path length considered through the Edenhofer et al. (2023) maps for estimating the dust complexity is therefore 1.05 kpc.
@@
After performing all the cuts, we are left with about 19,100 sightlines per posterior sample for this analysis. Those sightlines are shown at the bottom of Figure 5. The map shown is the sum of the binary masks over the 12 posterior samples. Note that the only discrepancies between the 12 posterior samples are at the edges of the selected regions. The rest of the map is either 12 or 0.
@@ 4.2 Dust Complexity
For this analysis, we experiment with various values for $\Delta_{n}$ and #$\#_{\rm voxels}$ and plot the resulting peaks identified for each variation. We find a range between $\Delta_{n}~{}=~{}1.94~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}3~{}\times~{}10^{-6}$ mag/pc) and $\Delta_{n}~{}=~{}3.87~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}9~{}\times~{}10^{-6}$ mag/pc) and #$\#_{\rm voxels}~{}=~{}3$ (21 pc) and a range between #$\#_{\rm voxels}~{}=~{}5$ (35 pc) recover the visually identified peaks. We report our results for the fiducial values of $\Delta_{n}~{}=~{}4.52~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}7~{}\times~{}10^{-6}$), and #$\#_{\rm voxels}~{}=~{}3$ (21 pc). However, we find our results to be robust to all the variations we test for within the ranges mentioned.
We run the dendrogram algorithm on each sightline separately, considering the entire sightline from 69 pc to 1.25 kpc. However, we only keep identified components with peaks that are radially farther than the distance at which the extinction in that sightline reaches 50 mmag (Section 4.1). If the distance of the peak of a component is farther than this threshold but part of the component is below that threshold (Figure 6), we still consider the part of the component that is below that threshold when integrating over the $n_{\rm H}$ of that component.
@@
We calculate Equation 10 for each sightline to quantify its dust complexity. For the sightline in Figure 6, for example, $\mathcal{N}_{c}$=2.16. We show a map of $\mathcal{N}_{c}$ for the sightlines we select in Section 4.1 at the top of Figure 7.
@@ 4.3 Nearest-Neighbor Matching
To examine the effect of line-of-sight dust complexity on polarization fraction, we compare sightlines that have the same total column densities but very different 3D dust distributions. We define low- and high-dust complexity sightlines as sightlines with $\mathcal{N}_{c}\leq 1.1$ and $\mathcal{N}_{c}\geq 1.5$, respectively. The goal is to compare sightlines with different dust distributions, so the particular $\mathcal{N}_{c}$ threshold values are less important. We start with these values for our fiducial analysis because they are similar to the thresholds used in Pelgrims et al. (2021). The Pelgrims et al. (2021) analysis used a Gaussian decomposition of 3D maps of the neutral hydrogen emission line, where the third dimension is radial velocity. They used $\mathcal{N}_{c}$=1 and $\mathcal{N}_{c}\geq 1.5$ in their analysis. However, for the data and mask we use in our analysis, we find that across the 12 posterior sample maps from Edenhofer et al. (2023), only 8-31 sightlines have $\mathcal{N}_{c}$=1, i.e., a single dust component. Since this does not represent a large enough sample size, we define the low-complexity bin as sightlines with $\mathcal{N}_{c}\leq 1.1$. This increases the number of sightlines in that bin to around 1,000 per posterior sample. However, we also confirm that our results are robust to variations in these thresholds (Section 4.6).
@@
We perform nearest-neighbor matching with no replacement between the sightlines in the low- and high-complexity bins. We pair up each low-complexity sightline to the high-complexity sightline with the closest $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ values based on the Manhattan distance using the ball-tree algorithm. For each matched pair, we subtract the $p_{353}$ of the sightline with the lower complexity from the $p_{353}$ of the sightline with the higher complexity. We take the average of the differences over the matched pairs, $\Delta p_{353}$, to test whether sightlines with higher dust complexity have higher depolarization levels on average than sightlines with lower dust complexity, i.e., $\Delta p_{353}<0$.
@@ 4.4 Statistical Tests
We determine the statistical significance of our results through permutation tests. We perform the analysis described in Section 4.3 and obtain $\Delta p_{353}^{s}$ for each of the 12 posterior samples $s$. We also perform null tests and obtain a distribution of $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample $s$. We calculate a p-value as the proportion of $\Delta p_{353}^{s,\,{\rm null}}$ that are equal to or more extreme than $\Delta p_{353}^{s}$.
For the null tests, instead of separating sightlines into high- and low-complexity bins, we randomly select 25% of sightlines (5,050 sightlines) to be in bin 1 and 25% to be in bin 2. We then run the same analysis on those 2 groups, pairing them up based on $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ and subtracting $p_{353}$ of the sightline in bin 1 from $p_{353}$ of the sightline in bin 2 for each pair. To ensure we are not biasing the null test by matching neighboring sightlines which have similar $p_{353}$, for each run, we randomly alternate between selecting the sightlines in bin 1 to be in the Northern Galactic hemisphere and the sightlines in bin 2 to be in the Southern Galactic hemisphere and the other way around.
@@
In addition to checking the statistical significance of $\Delta p_{353}^{s}$ for each of the 12 posterior samples, we also report the mean and standard deviation of $\Delta p_{353}^{s}$ across the posterior samples. The error propagated due to the uncertainty on $p_{353}$ is 2 orders of magnitude less than the result. The uncertainty is dominated by the variation in the result across the 12 samples.

### v2

@@ 3.2 Testing the Dependence of the Dust Polarization Fraction on the Magnetic Inclination Angle
Using 3D dust extinction maps and a model of the Local Bubble surface geometry, we can measure the angle between the line of sight and the local magnetic field orientation projected onto the Bubble’s surface in each direction. If the magnetic field lines were tangential to the Local Bubble surface as discussed in Section 3.1, the angle between the line of sight and the Bubble’s surface should on average be correlated with the measured dust polarization fraction along sightlines where the dust is concentrated in the Local Bubble wall. Therefore, by quantifying the correlation between this angle and the dust polarization fraction for different sightlines, we can test whether there is measurable evidence that the magnetic field traced by the dust polarization is preferentially tangential to the Local Bubble surface.
@@
$Q~{}=-\int\alpha\epsilon\rho\cos{2\psi}\cos^{2}{\gamma}\,ds,$ (3)
$U~{}=-\int\alpha\epsilon\rho\sin{2\psi}\cos^{2}{\gamma}\,ds,$ (4)
where $\rho$ is the volume density, $ds$ is a distance segment along the line of sight, $\epsilon$ is the dust emissivity, $\alpha$ is a coefficient defined in Equation 15 of Padoan et al. (2001) that is a product of polarization efficiency factors, such as the degree of dust alignment and the dust grain polarization cross-section, and $\psi$ is the angle between the projection of the magnetic field on the plane of the sky and South, and the Stokes parameters are given in the COSMO polarization convention. These equations show the dependence of the dust polarization fraction on $\gamma$$\psi$$\alpha$, $\rho$, and $\epsilon$ and their variations along the line of sight. At large angular scales and away from the Galactic plane, the variation in the observed dust polarization fraction is dominated by $\gamma$ and $\psi$ (Hensley et al., 2019; Planck Collaboration et al., 2020a).
@@ 3.3 Magnetic Fields in the Local Bubble Wall
    For the Local Bubble geometry, we use the 3D model developed by Pelgrims et al. (2020). Since this model is constructed based on the 3D dust map of Lallement et al. (2019), we use this dust map in this subsection. Example slices of the map with the model overlaid are shown at the top of Figure 2. The resolution of this dust map is 25 pc. We use a 3D Gaussian kernel with a standard deviation of 25 pc to smooth the data as done in Pelgrims et al. (2020). This smooths out spurious small-scale fluctuations in the data product that may affect the results.
@@
    $\alpha_{i}~{}=~{}\arccos{\left(\frac{\nabla A_{{V},i}^{\prime}\cdot\vec{r_{i}}}{|\nabla A_{{V},i}^{\prime}||\vec{r_{i}}|}\right)},\,\alpha_{i}~{}\in~{}[0,\pi/2],$ (5)
@@
Figure 5: A 2D histogram showing the joint distribution of the dust polarization fraction and $\theta$, the angle between the line of sight and the plane tangent to the surface of the Local Bubble (Equation 8) for the sightlines within the mask in Figure 3. There is no significant correlation between these two quantities.
@@
3.4 No Detected Imprint of the Local Bubble Wall on the Dust Polarization Fraction
@@
We do not find any correlation between $p_{353}$ and $\theta$ from Equation 8 over the mask defined in Figure 3. The joint distribution is shown as a 2D histogram in Figure 5. The Spearman rank coefficient, which is agnostic to the functional dependence between $p_{353}$ and $\theta$, is $3~{}\times~{}10^{-3}$. This lack in correlation persists for each of the masks described in Section 3.2. This indicates that at least one of the assumptions described in Section 3.1 is not valid. In other words, either the magnetic field is not generally tangent to the Local Bubble surface, the dust polarization is not dominated by the dust within 50 pc of the surface as defined by the Pelgrims et al. (2020), or a combination of these possibilities.
@@
We also test whether we find an anti-correlation between $\theta$ and the absolute value of the Faraday rotation measure (Section 2.3) but do not find any evidence for it. Therefore, we do not find an imprint of the detailed 3D geometry of the Local Bubble wall on the polarization statistics.
@@
O’Neill et al. (2024) provide a map of the inclination angle between the Local Bubble wall as defined by their model and the plane of the sky. We also test for a correlation between this angle and the dust polarization fraction. We use a similar sightline selection criterion, replacing the Lallement et al. (2019) map with the Edenhofer et al. (2023) map and the Pelgrims et al. (2020) model with the O’Neill et al. (2024) model. We calculate a Spearman rank coefficient of 0.04, i.e., we do not find any correlation.
4 Imprint of Dust Complexity on Dust Polarization Fraction
@@
We continue the investigation of how the 3D distribution of dust impacts the dust polarization fraction beyond the Local Bubble, taking into account the distribution of the dust in an extended volume around the Sun. We use the 3D dust maps constructed by Edenhofer et al. (2023) in this section. The benefits of these maps are that they have high resolution and extend radially up to 1.25 kpc away from the Sun.
@@
If contributions to the polarized dust emission originate from regions along the line of sight with differently oriented magnetic fields, the integrated signal will be depolarized relative to emission from a region with uniform magnetic fields. We postulate that sightlines with multiple dust components that are separated in distance and contribute similarly to the total column density are more likely to have substantial dust emission originating from regions with differently oriented magnetic fields. We test the hypothesis that on average, for sightlines with the same column density, the ones for which the dust is distributed into multiple components at different distances with similar contributions to the total column density are associated with higher levels of depolarization than those for which the dust contribution to the total column density is concentrated. Another way to state our hypothesis is that given two sightlines at the same column density, the one with a more complex 3D dust distribution will have a lower dust polarization fraction on average.
@@ 4.1 Sightline Selection
Edenhofer et al. (2023) use estimates of stellar distances and extinctions from Zhang et al. (2023) to construct their 3D dust maps. However, Edenhofer et al. (2023) found that their estimated extinction disagrees with Zhang et al. (2023) where the radially integrated differential extinction is below 50 mmag or above 4 mag. We therefore use those thresholds to mask sightlines where the dust differential extinction is likely to be significantly over- or under-estimated. The distance to those thresholds for a given sightline varies slightly for the 12 different posterior samples, so we apply a slightly different mask to each posterior sample. We only consider sightlines where the differential extinction integrated radially outwards reaches 50 mmag within 200 pc of the Sun. A threshold higher than 200 pc would include more sightlines in the selection, but it would shorten the minimum path length considered. Because in Section 4.5, we compare our results with the Leike et al. (2020) maps, which only extend out to 270 pc, we find 200 pc to be a good balance between having a large enough sample size and minimum path length. The minimum path length considered through the Edenhofer et al. (2023) maps for estimating the dust complexity is therefore 1.05 kpc.
@@
After performing all the cuts, we are left with about 19,100 sightlines per posterior sample ($\sim 40\%$ of the sky) for this analysis. Those sightlines are shown at the bottom of Figure 6. The map shown is the sum of the binary masks over the 12 posterior samples. Note that the only discrepancies between the 12 posterior samples are at the edges of the selected regions. The rest of the map is either 12 or 0.
@@ 4.2 Dust Complexity
For this analysis, we experiment with various values for $\Delta_{n}$ and #$\#_{\rm voxels}$ and plot the resulting peaks identified for each variation. We find a range between $\Delta_{n}~{}=~{}1.94~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}3~{}\times~{}10^{-6}$ mag/pc) and $\Delta_{n}~{}=~{}3.87~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}9~{}\times~{}10^{-6}$ mag/pc) and between #$\#_{\rm voxels}~{}=~{}3$ (21 pc) and #$\#_{\rm voxels}~{}=~{}5$ (35 pc) recovers the visually identified peaks. We report our results for the fiducial values of $\Delta_{n}~{}=~{}4.52~{}\times~{}10^{-3}~{}{\rm cm}^{-3}$ ($A_{V}^{\prime}~{}=~{}7~{}\times~{}10^{-6}$), and #$\#_{\rm voxels}~{}=~{}3$ (21 pc). However, we find our results to be robust to all the variations we test for within the ranges mentioned.
We run the dendrogram algorithm on each sightline separately, considering the entire sightline from 69 pc to 1.25 kpc. However, we only keep identified components with peaks that are radially farther than the distance at which the extinction in that sightline reaches 50 mmag (Section 4.1). If the distance of the peak of a component is farther than this threshold but part of the component is below that threshold (Figure 7), we still consider the part of the component that is below that threshold when integrating over the $n_{\rm H}$ of that component. We find that the peak with the highest dust column density lies within 270 pc from the Sun for most sightlines within our mask.
@@
We calculate Equation 10 for each sightline to quantify its dust complexity. For the sightline in Figure 7, for example, $\mathcal{N}_{c}$=2.16. We show a map of $\mathcal{N}_{c}$ for the sightlines we select in Section 4.1 for one of the 12 posterior samples of the Edenhofer et al. (2023) maps at the top of Figure 8.222We make the $\mathcal{N}_{c}$ maps for the 12 posterior samples of Edenhofer et al. (2023) publicly available at https://doi.org/10.7910/DVN/IW09AE (Halal et al., 2024)
@@ 4.3 Nearest-Neighbor Matching
To examine the effect of line-of-sight dust complexity on the dust polarization fraction, we compare sightlines that have the same total column densities but very different 3D dust distributions. We define low- and high-dust complexity sightlines as sightlines with $\mathcal{N}_{c}\leq 1.1$ and $\mathcal{N}_{c}\geq 1.5$, respectively. The goal is to compare sightlines with different dust distributions, so the particular $\mathcal{N}_{c}$ threshold values are less important. We start with these values for our fiducial analysis because they are similar to the thresholds used in Pelgrims et al. (2021). The Pelgrims et al. (2021) analysis used a Gaussian decomposition of 3D maps of the neutral hydrogen emission line, where the third dimension is radial velocity. They used $\mathcal{N}_{c}$=1 and $\mathcal{N}_{c}\geq 1.5$ in their analysis. However, for the data and mask we use in our analysis, we find that across the 12 posterior sample maps from Edenhofer et al. (2023), only 8-31 sightlines have $\mathcal{N}_{c}$=1, i.e., a single dust component. Since this does not represent a large enough sample size, we define the low-complexity bin as sightlines with $\mathcal{N}_{c}\leq 1.1$. This increases the number of sightlines in that bin to around 1,000 per posterior sample. The number of sightlines in the $\mathcal{N}_{c}\geq 1.5$ bin is around 13,000. However, we also confirm that our results are robust to variations in these thresholds (Section 4.6).
@@
We perform nearest-neighbor matching with no replacement between the sightlines in the low- and high-complexity bins. We pair up each low-complexity sightline to the high-complexity sightline with the closest $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ values based on the Manhattan distance using the ball-tree algorithm. The number of pairs is, therefore, equal to the number of sightlines in the low-complexity group, which is the smaller group. For each matched pair, we subtract the $p_{353}$ of the sightline with the lower complexity from the $p_{353}$ of the sightline with the higher complexity. We take the average of the differences over the matched pairs, $\Delta p_{353}$, to test whether sightlines with higher dust complexity have higher depolarization levels on average than sightlines with lower dust complexity, i.e., $\Delta p_{353}<0$.
@@ 4.4 Statistical Tests
We determine the statistical significance of our results through permutation tests. We perform the analysis described in Section 4.3 and obtain $\Delta p_{353}^{s}$ as the mean over the pairs for each of the 12 posterior samples $s$. We perform a permutation-based null test in which we randomly choose one slightline to subtract from the other in each pair, rather than always subtracting $p_{353}$ of the low-complexity sightline. We repeat this 10,000 times and obtain a distribution of $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample $s$. We calculate a p-value as the proportion of $\Delta p_{353}^{s,\,{\rm null}}$ that are equal to or more extreme than $\Delta p_{353}^{s}$.
We additionally use an alternative null test. Instead of separating sightlines into high- and low-complexity groups, we randomly select 25% of sightlines (5,050 sightlines) to be in one group and 25% to be in the other group. We then run the same analysis on those 2 groups, pairing them up based on $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ and subtracting $p_{353}$ of one group from that of the other for each pair. To ensure we are not biasing the null test by matching neighboring sightlines which may have similar $p_{353}$, for each run, we randomly alternate between selecting the sightlines in one group to be in the Northern Galactic hemisphere and the sightlines in the other group to be in the Southern Galactic hemisphere and the other way around. We find that sightlines in the Northern Galactic hemisphere tend to have higher $p_{353}$ on average than sightlines in the Southern Galactic hemisphere, so $\Delta p_{353}^{s,\,{\rm null}}$ will be biased towards positive or negative values based on which of the two groups is selected from which hemisphere. Using this version of the null test, therefore, takes this bias into account and yields a more conservative estimate of the significance of our hypothesis test. The permutation-based null test, however, guarantees the same number of pairs as the hypothesis test, so we use that version as the main null test when reporting the results. We find that our results are consistent regardless of which null test we use.
@@
In addition to checking the statistical significance of $\Delta p_{353}^{s}$ for each of the 12 posterior samples, we also report the mean and standard deviation of $\Delta p_{353}^{s}$ across the posterior samples. The error propagated due to the uncertainty on $p_{353}$ is 2 orders of magnitude less than the result. The uncertainty is dominated by the scatter from the 12 posterior samples of the 3D dust map.